OpenStudy (anonymous):

Simplify: 5 the square root of 3 (2 the square root of 3 + 6 the square root of 19) Last one @jim_thompson5910

OpenStudy (anonymous):

\[5\sqrt{3} (2\sqrt{3} + 6\sqrt{19})\]

jimthompson5910 (jim_thompson5910):

step 1 is to distribute

OpenStudy (anonymous):

What number? 5?

jimthompson5910 (jim_thompson5910):

the whole number \(\large 5\sqrt{3}\)

jimthompson5910 (jim_thompson5910):

you multiply that by EVERY term inside the parenthesis so you're multiplying \(\large 5\sqrt{3}\) and \(\large 2\sqrt{3}\) then you're multiplying \(\large 5\sqrt{3}\) and \(\large 6\sqrt{19}\)

jimthompson5910 (jim_thompson5910):

after all this, you add them together (combine like terms if possible, if not, then just leave it alone)

OpenStudy (anonymous):

10 sqrt(9) and 30 sqrt(57)

OpenStudy (anonymous):

Thank you.

jimthompson5910 (jim_thompson5910):

now you can simplify things a bit, so you're not completely done

jimthompson5910 (jim_thompson5910):

what is the square root of 9?

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

so 10*sqrt(9) = 10*3 = 30

jimthompson5910 (jim_thompson5910):

which means the final answer is \[\large 30 + 30\sqrt{57}\]